calcium is a group 2 metal which reacts with nitric acid according to the balanced symbol equation:

Ca + 2HNO3 - Ca (NO3)2 + H2

(relative atomic mass: H=1 N=14 0=16 Ca=40

0.2g of calcium metal was reacted with excess nitric acid. calculate the volume of hydrogen gas which was produced in this reaction


1
Expert's answer
2021-02-03T07:14:53-0500

Solution:

The balanced chemical equation:

Ca(s) + 2HNO3(aq) → Ca(NO3)2(aq) + H2(g)

According to the equation above: n(Ca) = n(H2)


Moles of Ca = Mass of Ca / Molar mass of Ca

The molar mass of Ca is 40 g mol-1.

Hence,

n(Ca) = 0.2 g / 40 g mol-1 = 0.005 mol


n(H2) = n(Ca) = 0.005 moles


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 0.005 mol of H2 occupies:

(0.005 mol × 22.4 L) / 1 mol = 0.112 L of H2 = 112 mL of H2

V(H2) = 0.112 L = 112 mL.


Answer: 112 mL of hydrogen gas (H2) was produced in this reaction.